an experimenter flips a coin 100 times and gets 54 heads. Test the claim that the coin is fair against the two-sided alternative.
Answer #1

Given that coin is flips 100 times and gets 54 heads.

x = 54 and n = 100

we have to test,

verses

Under Ho test statistic is,

So,

So test statistic is Z = 0.8

p-value is given by,

P-value = 2*P(Z >0.8) = 2*0.21 =0.42

So p-vlaue = 0.42 > 0.05 so do not reject null hypothesis

So We conclude that The coin is fair .